The following information concerns production in the Baking Department for December. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Dec. 1 Bal., 8,700 units, 4/5 completed 20,880
31 Direct materials, 156,600 units 281,880 302,760
31 Direct labor 79,040 381,800
31 Factory overhead 44,464 426,264
31 Goods finished, 158,700 units 412,272 13,992
31 Bal., ? units, 2/5 completed 13,992

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit $
2. Conversion cost per equivalent unit $
3. Cost of the beginning work in process completed during December $
4. Cost of units started and completed during December $
5. Cost of the ending work in process $

b. Assuming that the direct materials cost is the same for November and December, did the conversion cost per equivalent unit increase, decrease, or remain the same in December?
Answer #1

a1) Direct material cost per equivalent unit = 281880/156600 = 1.80

a2) Equivalent unit of conversion = (8700*1/5+150000+6600*2/5) = 154380

Conversion cost per equivalent unit = (79040+44464)/154380 = 0.80

a3) Cost of the beginning work in process completed during December = (8700*1/5*.80)+20880 = 22272

a4) Cost of unit started and completed = 2.6*150000 = 390000

a5) Cost of ending work in process = (6600*1.8+2640*.80) = 13992

b) Direct material cost in beginning work in process = 8700*1.8 = 15660

Conversion cost in beginning WIP = 20880-15660 = 5220

Conversion cost per equivalent unit in beginning WIP = 5220/6960 = 0.75

Conversion Cost increase by 0.05